OpenStudy (anonymous):

Determine if -1 is zero of the function above f(x)=-4x^3-5x^2-7x+2 @zzr0ck3r

OpenStudy (luigi0210):

Use synthetic division.

OpenStudy (anonymous):

@zzr0ck3r

OpenStudy (anonymous):

well i did my division @Luigi0210 and i didnt find -1 is zero

OpenStudy (anonymous):

i got it as not a rational function what would that mean would you be able to explain why exactly tho

OpenStudy (luigi0210):

Then it's not a zero~

OpenStudy (anonymous):

is that what you got?

OpenStudy (luigi0210):

The roots are: \[\LARGE x=0.23769, (-0.74384 \pm 1.2451i)\]

OpenStudy (anonymous):

so that means there's no zeros right

OpenStudy (anonymous):

@jim_thompson5910 @zzr0ck3r

OpenStudy (anonymous):

If you sub x=-1 into the x portion of the function, you would notice that the function does not yield 0, thus x=-1 is not a 0

OpenStudy (anonymous):

This is always the easiest way to check :)

jimthompson5910 (jim_thompson5910):

Plug in x = -1 and evaluate f(x) = -4x^3-5x^2-7x+2 f(-1) = -4(-1)^3-5(-1)^2-7(-1)+2 f(-1) = -4(-1)-5(1)-7(-1)+2 f(-1) = 4-5+7+2 f(-1) = 8 Since the result is NOT zero, this means that x = -1 is NOT a root (or zero) of f(x)

OpenStudy (anonymous):

explain how each theorem can be used in determining the solutions for f(x)=x^3+4x^2-x-10 fundemental theorem of algebra and descartes rule of signs and rational root theorem @jim_thompson5910